Please show step by step please. Answers are given, just need to know how to work them.

12) On a cold day, you take in 4.2 L of air into your lungs at a temperature of 0
Answer #1

12) when pressure is constant

V1/T1 = V2/T2...

4.2/273 = V2/(273+37)...
V2 = 4.2*310/273 = 4.8 L..

13) from P*V = n*R*T...

n = P*V/(R*T) = 101.3**4.2/(8.314*300)=0.17 moles...

here R is universal gas constant = 8.314 L kPa K^-1 mol^-1
Answer #2

The volume of air is proportional to the absolute temperature.
T1 = 273 degrees
T2= 310 degrees,
So V2/V1 = 310/273
V1 = 310 / 273 * 4.2 L
= 4.8

partB

just use ideal gas law:
pV = nRT
101300*(4.20/1000) = n*8.314*(27+273)
n = 0.17 mole

partB